Find dy/dx: 2^x+2^y = 2^(x+y) answer me this question with full steps and formulaes

Find dy/dx:
2^x+2^y = 2^(x+y) answer me this question with full steps and formulaes
 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
for dy/ dx
 
2^x ln2 + 2^y ln 2  dy/dx = 2^(x+y) ln 2 (1+ dy/dx)
 
2^x + 2^y dy/dx = 2^(x+y) (1+ dy/dx)
 
hope it helps